### v4

# “There Has To Be a Lot That We’re Missing”: Moderating AI-Generated Content on Reddit

Abstract.
Generative AI has begun to alter how we work, learn, communicate, and participate in online communities. How might our online communities be changed by generative AI? To start addressing this question, we focused on online community moderators’ experiences with AI-generated content (AIGC). We performed fifteen in-depth, semi-structured interviews with community moderators on the social sharing site Reddit to understand their attitudes towards AIGC and how their communities are responding. Our study finds communities are choosing to enact rules restricting use of AIGC for both ideological and practical reasons. We find that, despite the absence of foolproof tools for detecting AIGC, moderators were able to somewhat limit the disruption caused by this new phenomenon by working with their communities to clarify norms about AIGC use. However, moderators found enforcing AIGC restrictions challenging, and had to rely on time-intensive and inaccurate detection heuristics in their efforts. Our results highlight the importance of supporting community autonomy and self-determination in the face of this sudden technological change, and suggest potential design solutions that may help.
Online Communities, Reddit, AI-Generated Content, Generative AI, Moderation
ccs: Human-centered computing Empirical studies in collaborative and social computing
1. Introduction
Recent advances in generative AI (OpenAI22; openai2023gpt4; ramesh2021zeroshot) have led to the proliferation of AI-generated content (AIGC) across the internet (Thompson23Discovered), causing online communities to adjust their policies and practices. Generative AI use, whether disclosed or not, may alter the social dynamics of online communities. For example, AI-generated text with subtle inaccuracies may pose a challenge for knowledge-sharing communities, like Stack Exchange (makyen23), and those where members engage in cooperative work, such as Wikipedia (gertner_wikipedias_2023). In addition, the perceived inauthenticity of AIGC may be seen as disruptive to social communities where people come together for human connection, such as some Facebook groups (fiesler_ai_2024). This disruption adds to the existing challenges faced by online communities, such as coordinated disinformation campaigns, which generative AI may make cheaper, more effective, and better targeted (goldstein23).
As many communities rely on content moderators for protection (grimmelmann2015virtues), it is important to understand and document moderators’ firsthand experiences responding to AIGC. Effective moderation can help communities stay civil and aligned during times of disruptive change by clarifying and enforcing community norms (Kiene16; kiesler11). In this new age of generative AI, it is likely that existing moderation strategies for protecting online communities will need to adapt in order to remain effective.
This paper qualitatively investigates the experience of moderating AIGC in self-governing communities, in order to reflect on the impact of generative AI on online communities. Not all communities will see AIGC as an issue, but we center our investigation on those who do, in the hopes of illuminating ways that platform designers and community stewards may support community autonomy and self-determination in the face of sudden, dramatic technological change (Seering20). How to provide such support is a natural question for the HCI and Social Computing research communities, a central goal of which has been to encourage the healthy functioning of online communities as they strive to meet a variety of individual and collective needs (Cheng17; kiesler11; kraut11; Lampe10).
To this end, we conducted fifteen in-depth, semi-structured interviews with moderators of communities on Reddit. We focus on Reddit, a social sharing and news aggregation site, where we can study a collection of relatively independent self-governing communities. Reddit is an useful platform to study (proferes21) due to the large variety of types of communities it contains and its unique governance structure wherein individual communities (known as subreddits) set and enforce their own rules to supplement site-wide policy (Chandrasekharan18; FieslerEtal2018rrc). The focus on Reddit thus allows us to explore how different attitudes and practices towards AIGC are organically emerging across communities. We use a qualitative investigation to explore communities’ perspectives on what constitutes AIGC and what makes it problematic or challenging.
Our work explores the following research questions:
  • RQ1: What are the attitudes of volunteer community moderators towards AIGC use in their communities?
  • RQ2: What experiences have moderators’ had with AIGC in their communities?
We focus our investigation on large subreddits with rules governing the use of AIGC. Our exploration of RQ1 yields insight into our participants’ ideological and practical reasons for opposing AIGC use in their communities. We find that our participants’ definitions of AIGC, including what makes it problematic, vary across communities. Our participants acknowledge some potential benefits of AIGC, but overall say that it does not belong in their communities and poses several specific threats, such as lowering the quality of conversations, straining community relationships, and amplifying attacks on community members.
Our exploration of RQ2 reveals that participants largely feel that clarifying community norms about AIGC has presently allowed them to manage the threats that it poses, to the point where AIGC is now no more disruptive than any other problematic content. However, participants who moderate large subreddits and have experience with coordinated influence campaigns report that they do believe AIGC has made such campaigns more common. We note the challenge of enforcing rules about AIGC and detail the heuristics and technical tools that participants report using in these efforts.
We discuss the implications of these findings for members of online communities, community moderators, and platform designers. We explore the relationship between community values and attitudes towards AIGC and discuss design solutions that may help communities stay aligned as norms evolve. We also discuss the implications of moderators’ current approaches to detecting AIGC and explore design solutions that could aid with this task.
2. Related Work
Below we summarize three relevant bodies of work that motivate our study. We first explore research into the nature of online communities. This work demonstrates that online communities vary widely in both their reasons for forming and their values, which suggests that community attitudes towards AIGC will be context-dependent. Next, we review research on threats to healthy online communities which may be amplified by generative AI, such as deception, low engagement, and harassment. We conclude with research on the already-challenging nature of content moderation work in self-moderating online communities, showing that the difficulty of identifying AIGC may further strain moderators and their relationships with their communities.
2.1. Online Community Types and Values
Online communities form for different reasons, have different values, and thus may feel threatened by AIGC in different ways. Early work by Preece (preece_empathic_1999) coined the term empathic communities to refer to sites of both emotional and factual communication, where member participation is motivated by a desire for something that professionals cannot provide: firsthand accounts from others with similar lived experiences. We suspect that empathic communities may see AI-written responses as problematic and reject them for their lack of lived experience. Lave and Wegner provide another useful concept, the community of practice (lave1991situated), which represents a group of people united by a desire to improve their skills by doing, sharing, and receiving feedback from one-another. Since these communities deeply care about process, they may regard contributions as inauthentic if they are made using generative AI rather than their traditional practice. A final useful concept is Jenkins’ idea of knowledge communities in which “members work together to forge new knowledge often in realms where no traditional expertise exists,”(jenkins_introduction_2006). These communities value accurate knowledge and expertise, which may be threatened by generative AI’s current tendency to produce inaccurate “hallucinations” (augenstein2023factuality). Kraut and Resnick supplement this conceptual literature with practical considerations for platform designers (kraut11) hoping to encourage such communities through systems of trust and reputation (resnick2002trust). These systems of trust may need to evolve to account for a world where contributions made with AI are indistinguishable from those produced by expert community members.
Reddit has been a frequent topic of study for empirical investigations into the different types and values of online communities (proferes21). Reddit’s hierarchical governance strategy, in which a set of site-wide policies are supplemented by community-derived and enforced rules, allows it to host many independent communities with distinct purposes and values. This model of community self-moderation (Seering20) allows us to study evolving attitudes and practices as they emerge from affected populations. Prior work has analyzed surveys of community members (weld_what_2022), public subreddit rules (FieslerEtal2018rrc), and records of moderator actions (Chandrasekharan18), to identify the different values of communities across the platform. Findings from these studies emphasize the diverse set of values held by different subreddits. While they identify some commonly occurring rules and values, such as rules against harassment and the importance of trust, they emphasize that rules are largely context-dependent. These findings suggest that Reddit on the whole will not have a unified stance towards AIGC and motivates our inquiry into the specific perspectives of moderators from communities with rules governing its use.
2.2. Generative AI and Threats to Healthy Online Communities
The HCI literature has identified varied threats to online communities which might be amplified by the use of generative AI, such as deception, low engagement, and harassment. Past studies have shown online communities to be vulnerable to strategic influence operations, and recent speculative work has hypothesized that generative AI will increase the volume and believability of such efforts (augenstein2023factuality; goldstein23). Indeed, early empirical studies have found AIGC to be as or more deceptive than human-written content (park_ai_2024) and detected AIGC in both fake online reviews (Oak24) and spam content used for scams (diresta2024spammers). Other studies of intentional deception in online communities have focused on the capability of algorithmically controlled social media accounts, or social bots, to deceive unsuspecting users. (Ferrara2016; menczer_addressing_2023). While LLM-powered social bots have already been detected on Twitter (yang23), our work looks beyond social bots to additionally include situations where humans use generative AI tools to partially or completely craft their online contributions. We hypothesize that humans posting undisclosed AIGC may be regarded as a subtler form of deception, based on findings from AI-Mediated Communication (AIMC) studies into the effects of AI on interpersonal communication (hancock_aimc_2020) which show that online community members trust each other less (Jakesch19) and regard each other more negatively (rae_effects_2024) when they perceive that the other is using AI.
Another threat to online communities is that of declining participation, decreasing the value that a community provides to its members. The HCI and CSCW literature provides techniques for encouraging both contribution (kraut11) and commitment (ren_encouraging_2012) to online communities, such as establishing a strong group identity that allows strong ties to form between members. Studying how AIGC affects contribution and commitment in online communities is a relatively new area of research (wei2024understandingimpactaigenerated). Several studies of sites in the Stack Exchange network have demonstrated that the launch of ChatGPT caused an overall decline in website visits, question volume, and number of users, while increasing the complexity of questions asked (burtch_consequences_2024; sanatizadeh_exploring_2023; xue_can_2023). The authors attribute these effects to the relative ease and speed with which users can get simple questions answered by LLM-powered chatbots off-platform. Interestingly, the authors of one study (burtch_consequences_2024) contrast this effect with a measured null-effect on similar Reddit communities, which they suggest are able to avoid a decline in activity because of their emphasis on socialization, rather than pure information exchange. Recent work has also studied the impact of top-down bans on AIGC and found that banning the use of generative AI on Stack Exchange (wang_can_2024) led to a decrease in volume and quality of questions and answers. Our study builds on this line of inquiry by exploring whether or not the emergence of generative AI or community-imposed bans on the technology are threatening participation in self-governing communities organized around a variety of shared values and needs.
Finally, online communities are often sites of harassment and abuse. Better understanding online harassment and abuse is a frequent topic in the CSCW literature (cheng_anyone_2017; han_hate_2023), yet the impacts of AI-powered variants, such as malicious synthetic, or deepfake, media (farid_creating_2022; rini_deepfakes_2022) are much less studied. In response to what we see as a high potential for AI-powered abuse, our work explores how community moderators are thinking about these threats and what they are doing to try and deal with them.
2.3. The Challenges of Volunteer Community Moderators
Reddit communities who decide to enact rules governing the use of AIGC will turn to already-strained volunteer moderators to enforce these rules (li_all_2022; Wohn19). Moderation involves both “visible” and “invisible” work (Gilbert20), leading researchers to alternatively characterize it as “civic labor” (Matias19), “emotional labor” (Dosono19), and “care work” (gilbert23). The HCI community has a long history of addressing the needs of moderators (Jhaver19), which suggests that research may play a role in developing design solutions specific to AIGC. While existing research has focused on handling harmful content such as misinformation (Bozarth2023), the specific challenges of moderating AIGC have not been studied.
One perpetual challenge for moderators is maintaining alignment between themselves and the other members of their communities (Koshy23). Given the power dynamic between moderators and regular members, perceptions of unfair or biased moderation decisions can strain community relationships. Perceptions of bias are not unfounded: existing work has shown that individuals with marginalized identities are more likely to be the target of moderation actions (Haimson21; Thach24). We suspect that this will also be a problem when moderators must enforce rules about AIGC, as research on human perception of AIGC has found that humans cannot reliably distinguish between AI and human-generated text (clark21) or images (nightingale22), and use flawed heuristics to make such determinations (Jakesch_2023). Though little is known about AI detection in community self-moderation settings, recent work has shown that general determinations about AI use are biased according to race and gender stereotypes (mink24). Our study builds on this work by exploring the real-world AI detection practices of community moderators.
The CSCW research community has a long history of research on computational tools meant to aid moderators in their work (Chandrasekharan19; Jhaver19; Koshy23; Schluger22; Zhang20). One task where computational tools may help is detecting AIGC, although so far this task has proven challenging (yang_survey_2023). While commercial classifiers for AI-generated text claim to be effective (Crothers23), research on their practical application has found them to perform no better than random chance (weberwulff23), questioned their theoretical best-case performance (sadasivan2023), and even shown them to be biased towards classifying the text of non-native English writers as AI-generated (liang23). This research has called into question the feasibility of a reliable, general-purpose detection tool that will work on any snippet of text. Still, some studies have had success identifying AIGC in specific types of writing, such as news articles (hanley_machinemade_2024), conference peer reviews (liang_monitoring_2024), and crowd-worker output (veselovsky_prevalence_2023), suggesting that their techniques could potentially be adapted to other contexts. Our study will investigate how moderators think about these tools and currently incorporate them into their workflows.
3. Methods: Interview Study
To hear firsthand from Reddit moderators and to deeply explore a new area of research, we chose to conduct qualitative research via semi-structured interviews. We designed an initial interview protocol based on our research questions and updated it as themes emerged in early interviews. Participants were allowed to choose a medium that was comfortable to them for the interview: video call, audio call, email exchange, or message exchange via the Reddit platform. Regardless of channel, we posed our research questions one at a time and followed up in a semi-structured way. Pseudonymity is an important feature of the Reddit platform, so we did not ask for or record participants’ real names or offline identifiers. This study was IRB-approved. All participants read and accepted a consent agreement permitting their responses to be reproduced in research reports, along with the participants’ associated subreddits. To ensure that we reliably represented participants’ views, we contacted participants for feedback and approval on their included quotes before publication, following CSCW community recommendations (mcdonald19). We were able to solicit, receive, and integrate all participants’ feedback.
3.1. Crawl of Public Subreddit Rules
Our goal was to speak to moderators of communities who were actively governing the use of AIGC. To identify such communities, we launched a data collection effort to find subreddits with public community rules pertaining to AI use. We performed an automated crawl of the Reddit website in the summer of 2023 and extracted the text-based rules from all subreddits indexed by Reddit’s “Top Communities” list [^1] https://www.reddit.com/best/communities/1/, which includes public subreddits with more than ten subscribers. We crawled $n=337,399$ subreddits and found $n=87,596$ with non-empty rules and English as a primary language. We then performed a text search on the extracted rules to identify subreddits with rules related to AIGC. We acknowledge that this method is imprecise, but we found it sufficient for our exploratory analysis.
We found that 4% of all subreddits in the Top Communities index had rules about AI, but these rules were much more common in larger subreddits: when we restricted our analysis to top 1% of subreddits according to subscriber count, 20% had rules governing the use of AI. We thus decided to build our sample starting with the largest subreddits as these appear to be the most affected by AI at this moment. We also heard about moderating smaller communities from interviewees who moderated several communities of different sizes.
3.2. Recruitment
We recruited and interviewed participants during the summer of 2023. We manually examined the rules of subreddits in descending order of subscriber count. When we found a subreddit with either an explicit rule governing the use of AI or a rule that might implicitly cover AI use we added it to our sample, avoiding communities who banned research outreach. In addition, we also looked for public moderator discussions concerning AIGC in other forums, such as the r/ModSupport subreddit, and added those moderators’ subreddits to our sample. We then messaged the moderators of each subreddit in our sample via Reddit’s Modmail [^2] https://support.reddithelp.com/hc/en-us/articles/210896606-What-is-Modmail- feature, sending Modmail recruitment messages to 60 subreddits. We used two additional techniques to recruit participants beyond the largest subreddits. First, we conducted snowball sampling by asking moderators to forward our recruitment message to other potential participants. Second, we posted our recruitment call on several social media platforms. To ensure respondents were actual moderators, we asked for their Reddit username and the subreddit that they moderate, then verified that the users were indeed listed as moderators via their Reddit user profile. Participants were offered \$20 gift-cards as compensation. Overall, we interviewed fifteen participants who collectively moderated over 100 different subreddits with sizes ranging from less than ten members to more than 32 million (see participant information in Table 1).
3.3. Interviews
The first author conducted semi-structured 60 minute interviews with moderators about their experience with AIGC. The interviews used these key guiding questions:
  1. (1)
    What types of AIGC are moderators seeing in their communities?
  2. (2)
    What do moderators think are the motives of AIGC posters?
  3. (3)
    How are moderator and non-moderator community members responding to this new phenomenon?
  4. (4)
    What concerns do moderators have about AIGC and how severe are they compared to their other community concerns?
  5. (5)
    How do moderators identify AIGC?
  6. (6)
    What could help moderators address these concerns?
Interview transcripts were generated automatically from recordings and were manually reviewed and corrected by the first author. The first two authors then collaborated to iteratively code the interview content into categories, using an inductive thematic analysis method akin to Braun and Clarke’s Reflexive Thematic Analysis (braun21; Braun19). The two authors began by discussing themes that the first author observed while conducting the interviews and correcting the transcripts. From this conversation, an initial high-level set of codes emerged. Next, the first author open-coded all of the interview transcripts at the sentence-level, applying the initial codes and creating new ones where existing codes were insufficient. The two authors then reviewed the set of codes and refined them by merging similar codes and nesting those with a clear hierarchy. This round of coding produced no completely new themes, only new sub-codes, so the two authors agreed that the codes seemed fairly stable. Both authors then independently applied these codes to the same three interview transcripts and compared their results. Neither had created any new codes and the few cases of disagreement were resolved by further merging similar codes. This process yielded approximately 400 tagged phrases across all transcripts, which were grouped into the four high-level themes that we discuss below.
Table 1. Details for each interview. For each participant, we note the largest subreddit (according to subscriber count at the time of the interview) and the number of individual subreddits that they moderated.
Interviewee
Largest Moderated Subreddit
(Subscriber Count)
Subreddits
Moderated
Interview
Medium
Primary Threat
Concern
Level
1 r/todayilearned (32M) 25 Audio Call Amplifying Attacks High
2 r/food (23M) 25 Reddit Message Lowering Quality: More Off-Topic High
3 r/explainlikeimfive (23M) 3 Email Lowering Quality: Less Accurate Information High
4 r/explainlikeimfive (22M) 2 Audio Call Lowering Quality: Less Accurate Information High
5 r/WritingPrompts (17M) 3 Video Call Lowering Quality: More Off-Topic High
6 r/politics (8.3M) 25 Reddit Message Straining Relationships Medium
7 r/CryptoCurrency (6.7M) 8 Reddit Message Amplifying Attacks Medium
8 r/itookapicture (5.1M) 2 Email Lowering Quality: More Off-Topic Medium
9 r/NintendoSwitch (5.0M) 23 Email Lowering Quality: More Off-Topic Medium
10 r/Fantasy (3.4M) 1 Email Lowering Quality: More Off-Topic Medium
11 r/changemyview (3.4M) 2 Video Call Lowering Quality: Less Human Connection Low
12 r/changemyview (3.4M) 2 Audio Call Lowering Quality: Less Human Connection Low
13 r/AskHistorians (1.8M) 4 Video Call Lowering Quality: Less Accurate Information Medium
14 r/AskHistorians (1.8M) 1 Reddit Message Lowering Quality: Less Accurate Information Medium
15 r/GCTrading (22.7K) 3 Audio Call Amplifying Attacks Low
4. Findings
We share findings about subreddit moderators’ perspectives towards AIGC based on their personal moderation experiences, which center on four key themes: their general attitudes towards AIGC, the specific threats that they perceive AIGC poses to their communities, the ways that their communities are responding to the presence of AIGC, and the challenges of enforcing rules about AIGC.
4.1. Findings: General Attitudes Towards AIGC (RQ1)
While our participants acknowledged some potential benefits of AIGC, they reported generally opposing its use in their communities. Below we share insights into our participants’ various definitions of AIGC, then follow with details about their opposition to its use in their communities, and conclude this section with some potential benefits of AIGC that our participants also acknowledged.
4.1.1. Definitions of AIGC
Our interviews did not reveal a single consensus about the definition of “AIGC”, but we heard our participants use the term to refer to three different categories of content: images produced using a generative AI tool, text produced either entirely or partially with a generative AI tool, and posts made by automated accounts, or bots. Communities sometimes banned one of these categories of content while still allowing others.
We also heard different perspectives about what AIGC was problematic. Some participants categorically rejected AIGC, like a moderator of r/WritingPrompts [^3] To provide additional context, we introduce quotes with the name of the subreddit that the participant moderates. In the case where the participant moderates several subreddits, we include either the subreddit that the quote is referring to or, for more general quotes, the largest subreddit that the participant moderates. , who countered a community member’s argument that, “ChatGPT is another writing tool that authors are using and banning ChatGPT will be the same as banning a spellchecker,” with this response: “that argument is not good, because the level of effort required is clearly shifting quite dramatically.” Other participants acknowledged that AI-use can fall along a spectrum, from being used for minor editing to being used to completely write a post, and thus specific instances of AIGC are evaluated based on subjective characteristics. One participant who moderates r/itookapicture mentioned the importance of their perception of a poster’s [^4] We use the term poster to refer to both those posting top-level subreddit submissions as well as those posting comments in response to submissions. motive: “I don’t mind people using AI to better state their thoughts nor create art as long as it is not being used deceptively.” We also heard from participants who judged AIGC based on whether or not they deemed it to be the poster’s original work. For some participants, this was a matter of the degree to which AIGC still contained the poster’s own voice, while to others it was about AIGC being a poster’s “own intellectual contribution or content”:
r/AskHistorians moderator: So say, for example, we have an expert…perhaps they’re an expert on German history, but they don’t speak English all that well. So they write their answer in German, and then use ChatGPT to try to translate it…it is their own intellectual contribution or content, they’ve created this themselves, but then they just use this as a tool in a perfectly viable way.
The perceived intent and degree to which AIGC represents a poster’s own voice are subjective characteristics which could easily lead to disagreement about whether or not a specific example of AIGC is acceptable, which we revisit in later sections.
4.1.2. Opposition To AIGC
We dug deeper into the reasons behind our participants’ different notions of what makes AIGC problematic and found that our participants opposed AIGC in their communities for either ideological or practical reasons. Our participants voiced three different levels of objection to AIGC, which ranged from general to specific. Below we provide examples of the three levels of objection in the order of increasing specificity, starting with general objections to AI use per se, following with general objections to the use of specific AI tools, and concluding with objections to specific types of AIGC.
AIGC As Incompatible With Community Purpose
Ten participants reported opposing AIGC because they felt that using AI went against the purpose of their communities, which was to share posts created by humans. These participants objected to AI use in general, which suggests that there is no type of AIGC that they would welcome in their communities. These moderator’s opposition was ideological in nature, with one participant, a moderator of r/WritingPrompts, going so far as to say that they believed “[AIGC] is antithetical to everything that [our] subreddit is about”. These participants didn’t share objections to the quality of AIGC, but rather to the process through which it is created:
r/explainlikeimfive moderator: In our case, we do want someone to write these explanations. And we don’t want it done by some AI that has no clue what it’s talking about. And even if it does give an accurate answer, the purpose of this site is for people to write in their own words … there’s already Google, there’s already Wikipedia.
This moderator suggests that it is the personal, human-written nature of content in their community that differentiates them from other information sources like Google and Wikipedia. This quote suggests that the participant believes that contributors cannot “write in their own words” if they use AI, which we see as a context-specific norm that will likely differ between communities.
Current Generative AI Models Violate Community Values
Seven participants voiced ideological objections not to AI use per se, but rather to the use of current generative AI models, and by extension the content they produce, because they felt that certain characteristics of these models were incompatible with their communities’ values. While these characteristics of generative AI models may be difficult or unlikely to change, they are not inherent to the idea of AI, which suggests that these participants might not oppose AIGC in their communities were it produced by future generations of AI tools without these problematic characteristics. For example, one moderator of a creative subreddit shared that they object on ethical principles to AIGC because of the way that generative AI models are trained:
r/¡removed¿ [^5] We use ¡removed¿ to disassociate subreddits from participant quotes either at the participant’s request or when we find the quote together with the subreddit name potentially too revealing. moderator: Perhaps most importantly, as stewards of a creative space, we feel it’s our duty to support the real human artists, authors and other creatives whose work has been exploited to train the vast majority of these models without their knowledge and without credit or compensation.
This isn’t an objection to the use of AI per se, but rather to the use of AI tools that are trained according to the current paradigm, which the moderator feels is exploitative. This suggests that if model training methods evolve and address this concern then the moderator may no longer oppose AIGC in their community.
Another example is moderators’ stated opposition to generative AI models’ inability to reason, a trait which led a moderator of r/explainlikeimfive to describe ChatGPT as a “bullshit generator”. This is another objection based on process and not the content itself: even in instances where AIGC happens to be accurate, it is seen as problematic because it is produced by tools that work stochastically, as opposed to through reasoning. Some participants reported believing that generative AI models’ lack of relationship with the truth made them incompatible with the values of their communities:
r/AskHistorians moderator: History and the work thereof is something that I’d say AI will continue to find impossible for the foreseeable future, until such time as it can gain a sense of truth, or at least coherence.
Once again, this is not an objection to AI use per-se, but to the current paradigm of probabilistic generative AI models that are prone to hallucination. At this moment there may be no AI alternatives that would satisfy this objection, but this attitude does not close off future possibilities.
AIGC As Low Quality Content
Some participants didn’t object to AIGC in their communities on ideological grounds, but instead took a more practical stance and objected to AIGC because they saw it as low-quality content. While these were general attitudes, this opposition was often related to specific threats that they perceived from AIGC, which we will explore in greater detail in 4.2.
We heard from nine moderators who reported that AIGC did not meet the quality standards of their subreddits, like a moderator of r/AskHistorians who shared that AIGC, “tries to meet the substance and depth of a typical post…however, there are frequent glaring errors in both style and content.” Other moderators put it more bluntly, like a moderator of r/CryptoCurrency who noted that, “[AIGC] is not enjoyable to read”, or a moderator of r/Eldenring who shared: “[AIGC] is content that provides no value/discussion at all, it’s low effort content basically.” These participants did not take an ideological stance against AIGC, but evaluated it practically and saw it to add little value to their communities.
Other participants made practical objections to AIGC in their communities because they observed it being used in bad faith. Five participants, who moderated the largest subreddits, objected to AIGC because they saw it as a technique to rapidly accumulate Reddit karma [^6] https://support.reddithelp.com/hc/en-us/articles/204511829-What-is-karma, a practice known as karma-farming. Karma is a reputation points system that some communities use to limit who can make posts and karma-farming allows nefarious users to bypass this system. Some participants said they opposed AIGC because they had seen it used in the context of trolling, for example a moderator of r/politics shared that when they saw AIGC it was, “usually some form of trolling, or ‘outsourcing’ writing a comment the user wanted to make.” We also heard from participants who reported thinking that AIGC was primarily used to try and deceive other community members:
r/CryptoCurrency moderator: We’ve seen some ‘shills’ (aka companies or group of people) coming with AI content to promote their product…And of course promoting is detrimental, as it is often disguised as a ‘research’ post about that specific coin (for example), making it seem like a legit user researching a subject and that may influence buyers in comparison to a regular ad.
As we’ve just seen, our participants had a lot of opposition to AIGC use in their communities, but as we see next, some saw legitimate intentions behind its use.
4.1.3. Potential Benefits of AIGC
While the overall attitude that we heard from our participants towards AIGC in their communities was one of opposition, several participants did mention certain ways that AIGC could benefit their communities by increasing engagement. Participants’ attitudes towards the potential benefits of AIGC often came from what they perceived to be a legitimate, or at least good-faith, motivation behind AIGC posters. Five of our participants believed that a large amount of AIGC posting is motivated by a good-natured desire to participate. In particular, we heard from moderators who reporting thinking that community members were using AIGC to help complement their knowledge and writing ability:
r/AskHistorians moderator: We’re a very popular subreddit, and posts tend to be held up as examples of excellent content on Reddit. Some people may want to be a part of it, but not have the specific knowledge or resources to be able to contribute ‘properly’. Hence they turn to AI tools, hoping to provide a decent answer.
This participant was sympathetic with this motivation and engaged with the possibility that AIGC could be beneficial in certain situations.
In total we heard three separate moderators mention the potential for non-native English-speaking posters to use AI to improve their prose. These three moderators saw this as a legitimate desire to participate, but they emphasized that they wanted to hear the poster’s own voice, even if it was not perfect. We heard such a sentiment from a moderator of r/changemyview who shared: “I think AI as a tool is going to be useful…it’s useful for people who are poor writers, useful for people who aren’t strong English speakers. And when they need help, AI can fill in those gaps,” though they noted that it was important that such a contribution was still “sufficiently their words”. Of course, the definition of “sufficient” will be context dependent and may likely change over time as norms around AI use become more established.
One last attitude that we heard was that AIGC could be used to increase engagement in an otherwise inactive conversation. For example, one moderator speculated about the potential benefits of AIGC for platforms and provided a hypothetical example for us:
r/¡removed¿ moderator: So a subreddit that needs a little bit of life breathed into it, [Reddit] deploys some AI there and make a variety of posts. And then people comment and they see a whole bunch of ads. And everyone’s happy, right? A weekly automated thread isn’t necessarily a bad thing.
This moderator was not saying that their community had done this or that it aligned with their values, but rather that they believed this approach could be useful in certain contexts.
Although our participants were aware of communities elsewhere on Reddit who believed that the potential benefits of AIGC outweighed the negatives, none of our participants shared this perspective. Partially due to our sampling strategy, our participants reported that their general attitude of opposition to AIGC in their communities overshadowed any potential benefits.
4.2. Findings: The Specific Perceived Threats of AIGC (RQ1)
Our participants’ general attitudes towards AIGC were often rooted in more specific threats that our participants believe AIGC poses to their communities. These perceived threats included lowering the quality of posts, straining relationships between community members, and amplifying coordinated attacks. We provide more detail about each of these threats below.
First, it is worth noting that our participants had different perceptions of the severity of the threats from AIGC, with some seeing it as severe while others regarded it as something that could easily be managed through moderation. When asked if AIGC was one of their top concerns, a moderator for r/CryptoCurrency noted that, “Early 2023, I would’ve said yes, but right now it has calmed down quite a bit.” On the other hand, some participants did regard AIGC as one of their highest concerns:
r/explainlikeimfive moderator: It’s not our most common removal, by far. But personally, I would rate it as the most threatening concern…It’s often hard to detect and we do see it as very disruptive to the actual running of the site.
We see these two moderators as emblematic of a trend where our most concerned participants tended to moderate communities focused on authentic human connection, like r/explainlikeimfive, and the least concerned tended to moderate large subreddits that were less personal, such as r/CryptoCurrency. We will explore some of the detection challenges that the r/explainlikeimfive moderator references later in 4.4, but first we dig deeper into the why behind our participants’ different perceptions of the threat of AIGC to their communities.
4.2.1. Lowering Quality
Eleven moderators felt that posters using AIGC are “missing the point” of their communities. Subreddits exist for a wide range of reasons and adopt different values to meet their specific communities’ needs. We found that our participants assessed the quality of a contribution by considering how much it aligned with the values of their community. Our participants shared several specific ways that they thought AIGC lowered the quality of content in their communities, such as by decreasing the opportunities for authentic human connection, decreasing the reliability of information, and making conversations in the community more off-topic. All of these participants felt that low-quality content decreased the value that the community provides to its members and worried that it could result in decreased participation.
Less Human Connection
Seven participants mentioned the inherently inhuman nature of AIGC and expressed concern that this would lead to less authentic human connection in their communities. A moderator from r/changemyview summarized this sentiment well when they shared, “How can we change your view when you haven’t actually stated your view at all? This is ChatGPT’s view. And if I wanted to, I could go argue with ChatGPT. But that’s not helping here.” These participants believed that people come to their communities because they value connecting with other humans through dialogue and that AIGC does not meaningfully contribute to such conversations.
Less Accurate Information
Four participants who moderate communities that emphasize the accuracy of information (r/AskHistorians and r/explainlikeimfive) saw AIGC as threatening to their communities because AIGC is not necessarily factually sound. A moderator from r/AskHistorians shared that, “[inaccurate AIGC] undermines the whole intent of the sub, which is meant to give people that don’t have access…expertise in readable formats, from experts. And so, in the long term, it can undermine the trust that people have [in the community].” This participant described AIGC as “truthy-seeming” because of its tendency to include fake citations and felt that the appearance of being thoroughly researched increased the chances that readers would believe it.
More Off-Topic
Eleven participants said that they regarded AIGC as off-topic and felt that it would decrease the value of the community to its members. We heard several different reasons why our participants thought AIGC was off-topic. For example, six participants from communities that encourage or celebrate specific practices (e.g r/itookapicture, which is devoted to photography) said that they felt AIGC is off-topic because it is created without using their practice. When assessing the quality of a contribution to their community, these moderators were just as interested in the process used to create the contribution, as the contribution itself. The essence of their concern is captured by a moderator from r/WritingPrompts who noted, “[AIGC] is antithetical to everything the subreddit is about, which is about giving writers a chance to practice their craft and practice their skill.”
We also heard from five participants who thought that only original content was on-topic in their communities and that AIGC was not original content. For example, a participant who moderated r/WoW (a World of Warcraft subreddit where members frequently post fan art), said that, “On the art/other content side, we (on r/wow) consider it a matter of stealing w/o consent.” A moderator from r/AskHistorians echoed a similar sentiment when they described posting AIGC as, “taking words from something else. Even if it’s not a human, they are not your own words, and you are representing them as though they’re your own. So that’s plagiarism, right?”
4.2.2. Straining Relationships
In addition to feeling that AIGC was low-quality content, ten participants were concerned that AIGC would harm relationships in their communities. A moderator of a political subreddit was concerned about the general sentiment that emerged when legitimate users were falsely accused of posting AIGC. They noted that, “the possibility [of AIGC] fuels Redditors accusing other comments of being written by AI (which is a form of incivility).” A moderator of r/Zelda shared that the controversial issue of AIGC led to frequent off-topic arguments in their community. They reported that their main concern as a moderator was, “Do members feel safe and respected in our community,” and thus it was their their responsibility to remove, “arguments that shift away from civil debate towards petty fights / insults / trolling.” The participant saw that posts with AIGC often led to such arguments:
r/Zelda moderator: Many comment sections would fill with the same discussion points about the controversies of AI art…We would often see comments that attacked people for using AIGC, as well as attacks on people for gatekeeping art.
For this participant, banning AIGC from their community was thus a way to keep conversation civil.
4.2.3. Amplifying Attacks
Several participants mentioned the threat of AIGC being used to amplify malicious behavior like harassment and deception campaigns. However, moderators generally felt that existing techniques for countering these problems would be sufficient. For example, one participant shared concerns about AIGC being used as a tool for trolling, but they noted that they expected existing community defenses to be sufficient:
r/changemyview moderator: The one thing that I am worried about AI is that it does make trolling easier. You can generate larger amounts of trolling text and make it seem reasonable or seem like somebody wrote it for a lot longer…But we don’t tolerate trolling and actually it’s not that hard to tell when somebody’s trolling on our subreddit anyway. So it makes trolling easier [to do], but it’s something that I don’t think we’re going to be too stressed handling.
Another example is a moderator who shared about their experience with a ChatGPT-powered bot attack that was producing spam, which they were able to resolve with the help of Reddit admins:
r/AskHistorians moderator: A few months ago, we had an instance where we were subject to a bot attack that was using ChatGPT…Then we ended up reporting it to the admins who were able to take care of it through whatever they do on their back end.
In this example, the platform admins’ existing bot defenses were able to solve the problem, which made the participant view this kind of attack as a low concern.
The one exception to this was a moderator from r/news, which is one of the largest subreddits and frequently contains political content. This participant had seen their community targeted by, “a highly sophisticated deployment, an attempt to sway a narrative by utilizing accounts that had had their karma bolstered by automated commenting,” and were highly concerned that AIGC would make influence operations more frequent by facilitating karma-farming, which we discussed in 4.1.2 as an example of a bad-faith use of AIGC that our participants opposed. Our participant noted that while generative AI was much better at generating karma-worthy comments then earlier technologies, karma-farming was a threat only to the largest subreddits, which provide the greatest opportunity to gain karma.
While our participants reported different levels of concern about the threat that AIGC posed to their communities, all of them shared some degree of concern. Next, we share the changes that communities made to respond to these concerns.
4.3. Findings: Online Communities’ Responses to AIGC (RQ2)
Our participants shared various ways in which their communities have responded to concerns about AIGC, such as clarifying community norms and enforcing community rules.
4.3.1. Clarifying Community Norms
We heard from most participants that establishing a community stance towards AIGC was often a collaborative and iterative process between community members and moderators. Often it was the moderators who started the discussion, but other times it was the community:
r/museum moderator: It was the users who protested. Mildly: it’s a fairly polite subreddit. They said, we don’t believe that this belongs here. And so we had a blanket ban on it for a little while. And then we slowly allowed it to be integrated. And then there was a counter protest [against the integration].
We heard from most participants that they tried to develop rules for AIGC in consultation with their wider community, not only moderators, to ensure that they were aligned with the community’s interests. One participant described how the community started off permissive towards AIGC before eventually voting to ban it altogether:
r/CryptoCurrency moderator: At first (late 2022 early 2023) we let people use AI to write partial content in their posts *if they mentioned it*. But lots of abuse, people not mentioning it, and just overall quality decreasing, so [the community] decided to “ban” AI use.
This participant went on to describe the community’s process for publicly discussing issues, surfacing proposals for solutions, and soliciting votes on proposals from all community members—a process that their community used here to enact a rule banning AIGC.
All participants mentioned eventually codifying community norms into public subreddit rules. Some communities used rules that explicitly banned AIGC or bots, while others used more general rules (such as rules against plagiarism) as effective AIGC bans. Participants shared mixed results about the efficacy of these rules for decreasing the use of AIGC. We heard from seven participants that the volume of AIGC had decreased in response to enacting new rules, including from a moderator of r/AskHistorians who said, “People are using [AIGC] a little bit less now that they know what the rules are. They know that people don’t necessarily want it.” We also heard a moderator of r/explainlikeimfive express some skepticism about the effectiveness of rules: “I think it may stop some of the average users (to the extent that they actually read the rules, which is a reddit-wide problem.) Bots obviously don’t read rules.” This quote highlights that explicit rules can guide the behavior of good-faith users who want to act in accordance with the community’s values, but may not deter bad actors, especially if the rules are difficult to enforce.
4.3.2. Enforcing Community Rules
All of our participants’ communities devised—and tried to enforce—rules about AIGC. However, enforcement of these rules was often a challenge. We heard from three participants that needing to enforce rules against AIGC had increased their moderation workload:
r/AskHistorians moderator: It’s not fun to deal with having to assess whether or not something is [AIGC]. Reading through it like that, it’s a lot of extra work, right? Because you have to read through it all the way. It’s so fast for people to input this kind of content and a lot slower for people to assess it.
This imbalance between the amount of effort it takes to produce AIGC and the amount of effort that it takes to read it, especially since it tends to be stylistically more verbose than the average Reddit post, puts moderators at a disadvantage. We also heard from moderators who were concerned about alienating their community through false accusations. They noted that it took extra time and effort to get their decisions right, like a moderator of r/explainlikeimfive who shared, “You don’t want to be banning someone for using GPT when they don’t actually use it. So we might have to watch someone for a bit of time.” This need to spend time watching a community member for AIGC use is an unwelcome additional demand on the time of already over-burdened community moderators.
However, moderators were not the only ones involved in the rule enforcement process. Once a policy was in place, our participants relied on a combination of flagging by users, moderator review, automated tools, and platform action to enforce their rules. We heard from seven moderators that their community members often played an active role in rule enforcement, like a moderator of r/news who noted, “The first line of defense is the users themselves.” We also heard about instances where the load on the moderators became overwhelming, such as during the ChatGPT-powered bot-network attack discussed in 4.2.3, where moderators enlisted the help of Reddit platform admins to help them enforce their rules against bots.
Four participants mentioned that in addition to manual review, they used technical tools provided by the platform to enforce their rules, such as automoderator [^7] https://www.reddit.com/wiki/automoderator/, a programmable bot that scans and performs actions on all new posts. We also heard about additional tools available to the Reddit platform admins which were available for help in some severe cases of coordinated attacks:
r/news moderator: There’s the stuff that’s publicly available to every user, then you have the stuff that’s available to Moderator accounts, within the subreddit that the moderators are working on, and then you have the admin stuff—which is much more sophisticated than what we have.
The primary challenge when enforcing rules about AIGC is the difficulty of detecting it. Next we will dig more deeply into the techniques that moderators reported using to aid them with detection.
4.4. Findings: Detecting AIGC (RQ2)
Our participants reported using both custom heuristics and technical solutions to detect AIGC in their communities. They tended to report that their detection techniques largely work now but project that they will cease to be effective as generative AI increases in sophistication:
r/news moderator: We see the obvious stuff and we prune the obvious stuff…but there has to be a lot that we’re missing. And I imagine that it’s gonna get more and more sophisticated over time.
Below we discuss the challenges of detecting AIGC and provide examples of the heuristics and technical tools that moderators use to aid their work.
4.4.1. Detection Heuristics
Participants shared heuristics based on a variety of signals that they believe help them identify AIGC. To avoid aiding posters who are trying to evade these heuristics, we do not list the specifics of what our participants look for, but they broadly fell into the following categories: recognizable language patterns in posted content, details of users’ accounts and behavior, deviation from a community’s style norms, and inaccurate information.
Content Signals
Twelve participants mentioned that certain signals in a post’s content, such as keywords, phrases, or distinct forms, may make moderators suspect that a post is AIGC. A moderator of r/Eldenring noted that AI-generated text is, “very repetitive i.e it tries to justify a point but instead of doing that, it ends up repeating the same point over and over to the point of ad nauseum.” A moderator of r/itookapicture shared that with AI-generated images, “there is often something that feels a bit unnatural…The images look like the average of something, rather than a unique individual with all of the flaws that come with that.”
User Signals
Six participants reported that they often consider the details of a poster’s account, including their past posting behavior, when they suspect that a post is AIGC. One red flag that alerted participants was a dramatic change in the style of a poster’s prose. For example, a moderator of r/changemyview shared that suspicious posts were, “often very different in terms of writing from what the user posted in response to other users questions or comments.”
We heard from four participants that signals used to detect more traditional bots may also be relevant when identifying AIGC, like a moderator of r/Zelda who noted that, “accounts used for spam…there are other indicators to find those (account age, history,…) Even before ChatGPT, we would find copy-pasted or markov-generated comments with these same account indicators.”
Deviation From Subbreddit Style Norms
Four participants reported using their knowledge of their community’s particular communication norms to identify posts that don’t fit as AIGC:
r/AskHistorians moderator: It feels very formal, and it feels very different from the normal kinds of comments. There’s definitely sort of a genre to answering a question on r/AskHistorians that people who have been around for a while tend to all follow, just because there are certain ways to communicate historical knowledge in an open online space to a group of non experts.
Inaccurate Information
Four participants reported using their domain knowledge to identify posts as false or incoherent, which to them suggests AIGC:
r/museum moderator: In r/museum…there is a Canon, so to speak, of material. So it’s not like there’s suddenly going to be a brand new, you know, 16th century French academic painter that nobody’s ever heard of. And so something that seems incongruent to that situation would obviously raise an eyebrow.
4.4.2. Technical Tools
Though they acknowledged their unreliability, most of our participants reported using a variety of technical tools in detection efforts, such as black-box detection tools provided by third parties, and community-developed solutions. Participants mentioned programming the detection heuristics referenced above into automated tools like automoderator, which then apply the heuristics to every new post and flag anything suspect. Six participants mentioned manually copying posts into third-party detection tools to see if they were AIGC. A moderator of r/writingprompts reported that they had tested, but never deployed, a homespun bot that would apply third-party detectors automatically to all new posts.
At the same time, a moderator of r/changemyview mentioned that because they are aware of the unreliability of third-party detection tools, they rarely defer to them absolutely: “I don’t necessarily trust the tool–that’s why we have several layers.…We always try to talk to the user…But still, I think the tool is useful.” Two participants mentioned that the tools were least effective on shorter pieces of text, like a moderator of r/writingprompts who noted that the detector they were using, “was very bad with short comments. If a comment was like 100 words or a really short story–every poem got flagged as AI. And like, No, it’s not. It’s just short.”
5. Discussion
Our qualitative investigation offers two main research contributions: (1) a description of volunteer community moderators attitudes towards AIGC use in their communities and (2) a description of the moderators’ and their online communities’ responses to AIGC and the challenges that they’ve encountered governing its use. We place our work in dialog with Seering’s call to the CSCW research community to center community self-moderation (Seering20) and specifically see our study as motivated by the second “guiding question” proposed by Seering: “What are the processes of context-sensitivity in online community moderation, and how might they be better supported?” We believe that allowing communities to make context-sensitive decisions about AIGC is central to their self-determination and should be an important issue for the CSCW research community. Below we discuss the implications of our findings and potential ways that our research may be able to aid online communities during this period of technological change.
5.1. Community Values Shape Attitudes Towards AIGC
Our study reveals a spectrum of attitudes towards AIGC that seems to vary with the purpose and size of communities. Our exploratory crawl of subreddit rules (3.1) found that 4% of all listed subreddits had some rule governing the use of AI, but among the largest 1% of subreddits this number increased to 20%. This trend is consistent with earlier empirical work that has studied rules on Reddit and found community popularity to be the primary predictor of whether or not a subreddit has rules (FieslerEtal2018rrc). Our conversations with moderators suggest that certain types of communities, such as empathic communities (preece_empathic_1999) and  knowledge communities (jenkins_introduction_2006) may feel the most ideologically opposed to AIGC. We also see some examples of  communities of practice (lave1991situated) objecting to AIGC, but it is also likely that new communities of practice will emerge specifically devoted to AI-reliant practices [^8] https://www.reddit.com/r/aiArt/. On the whole, our findings align with those of recent work on the impact of generative AI on online communities (burtch_consequences_2024; wei2024understandingimpactaigenerated), which suggests that the social value of online communities may make them more important than ever in the age of generative AI. Platform designers and community stewards should focus on ways to increase this social value.
It will be interesting to observe how community norms towards AIGC evolve over time. Our interviews and analysis started less than a year after the launch of ChatGPT. Our findings indicate that community norms about AIGC have been in flux since this launch and will likely continue to change. It is possible that as people become more familiar with generative AI they will come to see it as compatible with their values, no different from other automated composition tools like auto-complete or spellcheck, which is an argument that we mentioned hearing in 4.1.1. If communities are to remain relevant, moderators will need to stay abreast of their community members’ stance on this evolving technology; ultimately, if community member and moderator perspectives on AIGC use are not aligned, Reddit’s distributed moderation system allows users to “agree to disagree” by leaving and finding other communities that share their stance (grimmelmann2015virtues).
At the time of writing, some platforms have begun incorporating generative AI into their products (lunden_linkedin_2023), likely in an effort to increase engagement. This may be aligned with platforms’ profit motives, but our interviews suggest that this is likely not aligned with the will of all platform users. As we mention in 4.1.3, generative AI may bring certain benefits to online communities, but given the strong current objections that we’ve surfaced, any attempts by platforms to integrate generative AI should be done via a collaborative design process with the affected communities (Gilbert20; Matias16; Matias18).
5.2. The Challenges of Establishing and Enforcing Norms
Our findings relate to and reflect on research about the power of norms to shape behavior in online communities (chancellor18; Chandrasekharan18; kiesler11; Lessig99) Platform designers may be able to aid community moderators with the work of establishing and enforcing community norms regarding AIGC.
5.2.1. Clarifying Norms in a Moment of Flux
Our interviews show that self-governing online communities are already adapting to the phenomenon of AIGC by clarifying their norms around its use and codifying these into rules. Prior research has shown that it is common for online communities to respond to sudden changes by making rules (SeeringEtal2019mec) and that this can be an effective strategy for navigating challenging transitions (Kiene16). Moderators sometimes struggle to identify AIGC, which makes enforcing rules about it a challenge; even so, discussing and clarifying rules can help inform the behavior of well-intentioned community members. It is likely that as generative AI grows in popularity and sophistication communities will continue to need to adapt their positions and practices in response. Our findings give us hope that communities will be able to successfully adapt, though it will require substantial effort on the part of both moderators and community members.
Given that AIGC is an emergent phenomenon, community norms around its use are still in flux. Community members and admins may have different notions about whether its use is acceptable or not, and this misalignment (Koshy23) may result in legitimate users inadvertently running afoul of their communities. Online communities may be able to lessen the threat of misalignment in the community by making their norms more explicit and known to their members, which has been shown to be effective with other moderation challenges (matias2019preventing). This approach is supported by our participants’ reports that the amount of AIGC has decreased since they enacted their rules banning it, although these reports are based on the moderators’ perceptions, and there are no reliable ways to independently verify them.
The importance of communicating norms about AIGC use also has implications for platform designers who may devise additional ways to ensure well-meaning posters are aware of community rules. For example, a community’s rules could be displayed in the authoring UI, perhaps with a checkbox that users must mark to indicate that they have read and are abiding by the rules. This is similar to how privacy policies are commonly presented. However, prior work has shown that privacy policies are rarely read (obar_biggest_2020), which suggests that adding a checkbox to the posting experience might only have the effect of adding friction to the experience and decreasing post volume. Another option would be to display the rules as placeholder text in the UI element where posts are written, which a previous study suggested as an elegant way to remind users of norms at the time of posting (Kiene16). That being said, displaying rules in a creative way may not be enough, as past research has demonstrated that many posters to online communities simply do not read the community’s rules before posting (JhaverAppling19). Communities concerned about this could also consider more thorough “onboarding experiences” to ensure new members are familiar with community rules, such as an interview with an existing member. However this approach places an additional time burden on both the existing community and the new member and may deter potential members who are wary of such a commitment.
Some social platforms are currently experimenting with incorporating generative AI features into their authoring tools (lunden_linkedin_2023). We suspect that the low effort required to use embedded generative AI tools will increase the overall amount of AIGC on platforms that add them. In order to respect the autonomy of the communities that sustain them, platforms should give individual communities the ability to decide for themselves whether or not to enable platform-provided generative AI tools. If a community opts out, this decision could be communicated to posters via the UI, perhaps by displaying the tools in a disabled state with an informative tooltip. Communicating the community’s stance on AIGC in this way may have the added benefit of also discouraging posting AIGC produced by off-platform generative AI tools.
5.2.2. Equitably Enforcing AIGC Bans
Our findings detail some of the challenges that moderators currently experience enforcing AIGC restrictions and bans. These challenges impact both moderators and community members, and our findings may help offer potential solutions.
Even before having to enforce rules about AIGC, moderators were stretched thin trying to protect their communities from problematic content (e.g. mis- and disinformation) and behavior (e.g. harassment, incivility, and influence campaigns). With detection of AIGC being an unsolved technical problem (edwards23), moderators are left to design and use heuristics to enforce restrictions or bans. Our findings show that these heuristics can be time-intensive. Worse, the heuristics might also be flawed, or lead to biases as people’s intuitions about AIGC are often wrong in predictable ways (Jakesch_2023). Such consistent biases could disproportionately affect minorities and individuals seeking legitimate participation (liang23), such as non-native speakers, newcomers, and gender or racial minorities. It is concerning that most of our participants were confident in the efficacy of their heuristics and unaware of any potential bias. Our participants’ current reliance on potentially biased heuristics suggests the importance of both educating moderators about these biases and conducting further studies to evaluate the efficacy of detection heuristics in the community moderation context.
Platform designers could aid the work of moderators by building moderation tools that surface the relevant signals that moderators use for their detection heuristics. Earlier work has suggested tools that surface “synthesized social signals,” or “social signals computationally derived from an account’s history,” to help online community members evaluate the credibility of other accounts (Im20). Such a tool could be useful here if it additionally surfaced information about users’ interactions with the system (e.g. time spent editing a post) that would give moderators a sense of the effort that authors put into their posts, though this may be difficult to implement in a way that respects users’ privacy. Additional information about a poster’s past behavior and other community memberships may also be relevant, similar to how these are useful in assessing if a user is likely to post misinformation (Bozarth2023; diakopoulos_diamonds_2010).
Given the combination of concerns about AIGC and the lack of tools to detect it, we expect to see communities developing their own, non-technical, communication norms for demonstrating authenticity, similar to past developments such as timestamping and triforcing (bernstein14), practices used in anonymous and pseudonymous online communities such as Reddit. These techniques are a way to signal in-group membership by presenting something that only an authentic poster would be able to present. Platforms may also want to update their authoring tools to support any new community-specific signaling practices that users develop to signal their authenticity.
6. Limitations and Future Work
Our qualitative study is based on interviews with moderators of some of the largest subreddit communities, but, naturally, may not cover sentiments that generalize across the entire Reddit platform. Future work could perform a platform-wide quantitative study that could quantify trends and patterns across subreddits of different size and type. Further, since AIGC is an internet-scale phenomenon, future studies should explore how moderators are responding on other platforms besides Reddit. We recruited from popular subreddits in an attempt to target those that are currently most affected, but our participants who also moderate smaller subreddits shared that their experiences and concerns are relevant to communities of various sizes. While we took steps to verify the accurate representation of our participants’ views, several emphasized that they can only speak to their own experience and do not claim to speak on behalf of their entire communities. Because our recruitment strategy targeted subreddits with explicit rules about AIGC, we also likely omitted those communities that have fully embraced AIGC and thus have not created any rules against its use. Additional studies could include moderators from such communities for a more complete picture of the range of ways that online communities are responding to AIGC, and to further explore the relationship between a community’s values and its stance on AIGC.
Our participants mentioned heuristics that they use to detect AIGC, but we have not empirically verified their accuracy. Existing studies have demonstrated the inaccuracy of detection heuristics in a lab setting (Jakesch_2023), but further research could measure the accuracy and bias of the detection heuristics of “experts” who are working in a context that they are deeply familiar with, such as community content moderators.
7. Conclusion
We performed a qualitative investigation into volunteer community moderators on the social sharing site Reddit’s experiences with, and attitudes towards, AI-generated content use in their online communities. Rather than try to represent the range of moderators’ attitudes and experiences across all of Reddit, we focus our study on the moderators of communities who have decided to restrict the use of AIGC, in the hopes of understanding how the CSCW research community may support these efforts towards community self-determination. We find both ideological and practical objections to AIGC that make our participants feel AIGC does not belong in their communities, despite also hearing about its potential to make it easier for some community members to contribute. On the whole our participants reported concern that allowing AIGC would decrease the quality of conversation and interactions in their communities, but they felt that the current restrictions that they put in place were effective, for now, at protecting their communities. These restrictions primarily served to clarify and communicate the community norms around an emerging issue with generally unsettled norms, but the actual enforcement of these restrictions is imperfect.
As generative AI usage continues to grow, the CSCW community needs to continue to pay attention to how it affects the dynamics, practices, and health of online communities. There is much that the research community can do to develop tools and practices that may help community moderators guide their communities through this period of technological change. As a starting point, this paper explores the perspectives of moderators at this critical juncture, where, as one moderator commented, “there has to be a lot that we’re missing.”
[…]